Grade 12th passMagical Mathematics[Interesting Approach]

If for 0o , e[(sin2x + sin4x + sin6x + ……)loge2] satisfies the quadratic equation x2-9x+8=0.

Find the value of sinx-cosx

sinx +cosx

rohan kanojia
15 Years agoGrade 12th pass

1 Answer

SAGAR SINGH - IIT DELHI
ApprovedApproved Tutor Answer15 Years ago

Dear student,

e[(sin2x + sin4x + sin6x + ……)loge2]

On solving we get 2^tan2x

x2-9x+8=0

this gives x=8,1

2^tan2x=1 and 2^tan2x=8

On critically examining we get x=pie/3

sin60-cos60/(sin60+cos60)=root 3-1/root 3+1

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

exp {(sin2x + sin4x + sin6x + ...... ∞) loge2

.log 2
1 sin x
sin x
2 e
2
e − ⇒ tan 2 x
eloge 2
⇒ tan2 x 2 satisfy x2 – 9x + 8 = 0 ⇒ x = 1, 8
∴ tan2 x 2 = 1 and tan2 x 2 = 8
⇒ tan2x = 0 and tan2x = 3
⇒ x = nπ and tan2x =
2
3
tan ?
?
?
???
π
and x = nπ ±
3
π
Neglecting x = nπ as 0 < x <
2
π
⇒ x =
3
π
∈ ?
?
?
??
? π
2
0,

cos x sin x
cos x
+
=
2
3
2
1
2
1
+
=
1 3
1
+
×
3 1
3 1


=
2
3 −1

cos x sin x
cos x
+
=
2
3 −1